The sum of two numbers is 34 and the difference is 8 What are the numbers?
Answer by ankor@dixie-net.com(22740)

The sum of two numbers is 34
a + b = 34
and the difference is 8
a - b = 8
What are the numbers?
use elimination
a + b = 34
a - b = 8
-----------addition eliminates b, find a
2a + 0 = 42
a = 42/2
a = 21
:
Ill let you find b , check the solutions in both equations